OpenStudy (anonymous):

Which of these is a simplified form of the equation 7z + 2 = 8 + 2z + 1z? A. 10z=6 B. 10z=10 C. 4z=6 D. 4z=10 @robtobey

OpenStudy (dtan5457):

Add like terms, and isolate the variable z.

OpenStudy (dtan5457):

7z+2=8+2z+z 7z+2=8+3z Can you take it from here?

OpenStudy (anonymous):

no @dt

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

7z+2=8+3z Subtract the 2 and 3z from both sides..

OpenStudy (anonymous):

5 and 1

OpenStudy (dtan5457):

@iambatman Can you explain this better?

OpenStudy (anonymous):

7 - 2 ,, 3-2

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

When you have a problem as such: 4x+1=4-x Don't be afraid of the the variables, the idea is to isolate x (get it by itself), so our first step in this example is add x on both sides, 4x+1+x=4-x+x you see on the left side it gets eliminated, -x+x = 0, so we're left with 4x+x+1 = 4, now you want to move the 1 on the right side, by subtracting the 1 on both sides (what you do to one side, you must also do the other, remember this) 4x+x=4-1 4x+x = 5x 4-1 = 3 putting this together 5x=3 now you see the 5 is "stuck" on to the x, which means it's being multiplied, and what's the opposite operation of multiplying? Division, so we divide by 5 on both sides so we're left with x, because 5/5 = 1, 1*x = x So our solution is x = 3/5

OpenStudy (anonymous):

so whats the answer ? @iambatman

OpenStudy (anonymous):

...

OpenStudy (anonymous):

I think it's time I give you a break from OS as you are not putting any effort and just looking for the answers, I suggest you try your questions on your own, and when you come back, you will have questions about where you got stuck or something about the concept.

OpenStudy (dtan5457):

@iambatman Well put.